# Causal Graph Neural Networks for Healthcare

### v3

@@ 1 Introduction
In 2019, a widely deployed risk prediction algorithm for the allocation of healthcare resources systematically underestimated the severity of the disease for Black patients by relying on healthcare costs as a proxy for health needs, despite Black patients receiving less aggressive treatment even when experiencing an equivalent burden of disease, resulting in discrimination affecting millions of patients annually [1]. That same year, a diabetic retinopathy screening system that achieved 94% accuracy in one hospital dropped to 73% in another, having learnt site-specific correlations, perhaps scanner characteristics or local clinical protocols, rather than causal disease mechanisms [2]. These failures exemplify the triple crisis of AI in healthcare: models collapse under distribution shift, perpetuate discriminatory patterns embedded in historical data, and struggle to provide reliable mechanistic insight that clinicians can verify [3, 4, 5]. The paradox of pneumonia risk prediction starkly illustrates this brittleness: models learnt to classify asthma patients as lower risk because they received aggressive early treatment in training hospitals, which could lead to dangerous under-treatment when deployed elsewhere [6]. The question facing healthcare AI is no longer whether systems can achieve impressive accuracy in retrospective test sets, but whether they can survive clinical deployment across diverse institutions, populations, and practice patterns and maintain both effectiveness and equity.
@@
This Perspective examines causal graph neural networks, methods that integrate causal inference principles with graph-based deep learning architectures, as a methodological framework for transforming AI in healthcare from reactive pattern recognition to proactive causal medicine. Unlike previous reviews that addressed GNNs in healthcare [36, 37] or causal inference in medicine [16, 38] separately, we focus specifically on their integration, methods in which both graph structure and causal reasoning prove essential for clinical impact. We analyse how causal principles enhance GNN architectures for healthcare tasks, moving beyond correlation to mechanism, and examine current applications that demonstrate clinical utility across psychiatric diagnosis, cancer treatment, physiological monitoring, and drug discovery. We also identify fundamental challenges in scaling these methods to real-world deployment, including computational complexity, validation without ground truth, fairness under distribution shift, and regulatory considerations. Our analysis reveals both transformative promise and substantial barriers: while CIGNNs achieve superior generalisation across healthcare institutions, identify biologically plausible mechanisms, and maintain fairness properties under distribution shift, they face computational requirements precluding real-time application, difficulty validating causal claims from observational data, and gaps between methodological sophistication and clinical interpretability.
@@ 2.1 Why biological systems demand graph-structured causal models
Biological systems exhibit causal relationships organised as networks across multiple scales of organisation. At the molecular level, gene regulatory networks encode transcriptional control in which regulatory proteins causally influence the expression of target genes, with modern single-cell RNA sequencing that captures the expression profiles of 20,000+ genes simultaneously [30]. Protein-protein interaction networks represent physical binding relationships that mediate cellular signalling, with causal perturbations propagating through these networks to generate phenotypic responses [18]. At the tissue level, brain networks comprise hundreds of regions with functional and structural connectivity encoding causal influences between neural populations, where activity in one region can drive or modulate activity in connected regions through synaptic transmission [43]. At the population level, disease comorbidity networks represent causal relationships where one condition increases risk for others through shared pathological mechanisms, diabetes causally increasing cardiovascular disease risk through vascular damage, inflammation, and metabolic dysfunction [19]. This inherent graph structure makes neural network architectures that explicitly represent and reason about network topology the natural framework for biomedical machine learning.
@@ 2.2 From association to intervention to counterfactuals
Counterfactual reasoning operates at the highest level of Pearl’s hierarchy, addressing patient-specific questions: would this patient have developed complications if we had chosen alternative treatment? Reasoning of this kind is essential for precision medicine, treatment optimisation, and retrospective analysis of clinical decisions [15]. The calculation of counterfactuals requires a three-step process, including abduction (inferring unobserved variables, including patient-specific factors, from observed data), action (modifying the causal model to reflect alternative intervention), and prediction (computing outcomes under the modified model) [44]. The CLEAR framework implements counterfactual generation for graphs through variational autoencoders that learn distributions over causally plausible graph modifications, allowing the exploration of alternative biological configurations with mechanistic validity [54]. In cancer treatment planning, this enables simulation of how different therapeutic targets would affect tumour evolution, identifying vulnerabilities that might not be apparent from observational data alone.
@@ 3.1 Disentangling causal signal from confounding
The DisC (Debiasing via Disentangled Causal Substructure) framework operationalises this separation through a dual-encoder architecture [52]. Rather than treating graphs monolithically, DisC maintains parallel GNN encoders: one trained to extract causal features that consistently predict outcomes across different environments, another capturing environment-specific spurious patterns. The contrastive learning objective enforces the independence between these representations, ensuring that causal and spurious features remain disentangled. For psychiatric diagnosis, this architectural separation was proven to be immensely effective. Traditional GNNs identified widespread hyperconnectivity in the cerebellar regions of depressed patients, a finding inconsistent with the clinical understanding that depression primarily influences affective rather than motor circuits. CI-GNN’s causal decomposition correctly identified sparse cerebellar involvement while highlighting strong connections around the left rectus, a region previously identified as critical for antidepressant treatment response prediction [53]. Evaluated on the REST-meta-MDD dataset (n=1,604), this alignment with established neurobiology demonstrates that causal disentanglement provides mechanistically meaningful insights, not merely improved prediction.
@@
The CI-GNN (Causality-Inspired Graph Neural Network) advances this concept through Granger causality-inspired conditional mutual information, providing principled measures of causal strength for graph edges [53, 52]. The model identifies influential subgraphs that demonstrate genuine causal connections through disentangled subgraph-level representation learning. In autism spectrum disorder evaluations on the ABIDE dataset, CI-GNN revealed tight connections within the somatomotor network consistent with motor impairments characteristic of ASD, findings replicated across independent cohorts and suggesting genuine causal relationships rather than dataset-specific artefacts [53]. The clinical imperative was that clinicians needed to know which connections to target through neurostimulation, pharmacotherapy, or behavioural interventions, not merely that ”connectivity is abnormal.” The IGCL-GNN framework extends these principles through multi-objective optimisation that simultaneously enhances causal features and suppresses environment-specific spurious patterns, recognising that robust clinical deployment requires balancing predictive performance on current datasets with generalisation to unseen environments [59].
@@
Notably, these findings depend on the chosen parcellation resolution. CI-GNN operates on 116-node brain atlases where each node aggregates thousands of neurons; causal relationships at this mesoscale resolution reflect functional circuit organisation rather than individual neuronal connectivity, with direct implications for which neurostimulation or pharmacological interventions could target the identified causal pathways. Figure 3 illustrates these architectural principles applied to functional neuroimaging data. While the DisC and iVGAE frameworks were originally developed on synthetic benchmark datasets, their conceptual translation here to brain connectivity analysis could demonstrate how causal disentanglement and interventional prediction could address the specific confounding challenges in clinical neuroimaging, separating genuine neural circuit dysfunction from global arousal states, scanner artefacts, or site-specific acquisition protocols.
@@ 3.2 Predicting interventional effects without experimental data
Precision medicine requires predicting patient-specific treatment responses, queries about interventions that often lack experimental data. The RC-Explainer (Reinforced Causal Explainer) uses reinforcement learning to discover which edges of the graph to intervene in to maximise the causal effect on the outcomes [60]. Rather than computing interventions symbolically, it learns a policy for selecting interventions, with each candidate evaluated not by correlation with outcomes but by causal impact computed through careful accounting for confounding and interaction effects. In drug combination discovery, this approach identified synergistic therapies in which individual drugs showed minimal effect but together produced substantial therapeutic benefit, combinations that were missed by correlation-based methods that cannot distinguish synergy from additive effects [60].
The CaT-GNN (Causal Temporal Graph Neural Network) implements interventional reasoning through architectural modifications encoding backdoor adjustment [61]. The framework first identifies which nodes represent causal factors versus environmental confounders, and then applies mixup augmentation specifically to environmental nodes. This selective augmentation breaks spurious correlations and preserves causal relationships, achieving robust prediction across different patient populations and clinical settings. For blood pressure monitoring, the CiGNN framework demonstrates the clinical utility of interventional prediction by first using the neural-accelerated causal discovery (FCI algorithm) to identify physiological signals causally related to blood pressure variation, then constructing spatiotemporal GNN architectures guided by the discovered causal structure [51, 62]. This causal feature selection reduces dimensionality from 222 initial wearable features to a sparse set of direct and indirect causal indicators while improving both accuracy and interpretability, identifying, for example, that specific PPG morphological features (e.g., pulse arrival time, waveform characteristics) causally influence blood pressure through arterial stiffness mechanisms.
@@ 3.3 Generating counterfactuals for personalised medicine
Counterfactual reasoning addresses the highest level of Pearl’s causal hierarchy: what would have happened to this patient under alternative treatment? Such patient-specific predictions are essential for precision medicine, treatment optimisation, and retrospective analysis of clinical decisions [15]. The CLEAR (Counterfactual Learning on Graphs) framework approaches counterfactual generation through variational autoencoders that learn distributions over possible graph modifications [54]. Rather than generating arbitrary alternative graphs, CLEAR learns which modifications are causally plausible, which edges could change, and it still maintains biological validity. In cancer treatment planning, this enables exploration of how different therapeutic targets would affect tumour behaviour, identifying unexpected vulnerabilities invisible in observational data alone.
The fundamental challenge in counterfactual reasoning is that interventions can create graph configurations far outside the training distribution [63]. When simulating disruption of a critical hub protein, we query a configuration likely never appearing naturally, and thus unrepresented in training data. Traditional neural networks fail catastrophically under such a distribution shift. The CSA (Causal Supervision for Attention) framework tried to solve this through attention-based counterfactual analysis [64]. Rather than generating explicit counterfactual graphs, CSA learns attention patterns that can be smoothly modified to simulate interventions, ensuring counterfactual predictions remain within the model’s reliable operating regime while exploring meaningful alternative scenarios. For epileptic seizure prediction, this enables in silico testing of interventions, simulating how disrupting specific neural circuits would affect seizure propagation patterns, guiding neurosurgical planning or neurostimulation protocols [65].
The CXGNN (Causal Explainer based on GNN-Neural Causal Models) framework provides a rigorous quantification of counterfactual effects through neural implementations of structural causal models [66]. By parameterising each structural equation as a neural network, CXGNN learns arbitrary non-linear causal relationships and stays true to the formal guarantees of SCM-based counterfactual computation. For mild cognitive impairment progression, CXGNN identified that hypertension, arrhythmia, and sex emerged as critical causal factors, with notable sex-specific differences: female patients showed stronger causal links between vascular factors and cognitive decline, while male patients exhibited more pronounced metabolic effects [66]. These mechanistic insights enable personalised risk assessment and targeted early intervention strategies. These counterfactual generation methods form the computational engine of Causal Digital Twins; the ability to simulate patient-specific responses to interventions never administered in reality constitutes the core functionality enabling in silico clinical experimentation.
@@ 3.4 Achieving robustness across heterogeneous clinical settings
While disentanglement methods (Section 3.1) achieve robustness by explicitly separating causal from spurious representations, invariant learning approaches pursue the same objective through a complementary mechanism: optimisation constraints that enforce feature stability across environments without requiring explicit architectural separation. Disentanglement provides interpretable causal-spurious decomposition; invariant learning provides deployment guarantees with potentially simpler architectures. Some methods, such as IGCL-GNN, integrate both principles.
@@
The IGCL-GNN framework operationalises invariant causal learning through multi-objective optimisation that enhances causal features and suppresses environment-specific spurious patterns [59]. The key insight is that robust clinical deployment requires balancing competing objectives such as maximising predictive performance on current datasets and also ensuring generalisation to unseen environments. In multi-site brain imaging studies, CI-GNN achieved mean accuracy of 0.71–0.73 across 17 acquisition sites compared to 0.64–0.69 for baseline methods, with consistent improvements in 13 of 17 sites despite individual site accuracy ranging from 0.60 to 0.85, demonstrating the value of learning causal invariances rather than site-specific correlations [52, 53]. This stability under distribution shift is not just a technical achievement, but a prerequisite for clinical trust—models that collapse when moved between hospitals cannot be deployed safely.
The CRec (Conditional Causal Representation Learner) framework addresses the robustness of drug recommendation through disentanglement of therapeutic effects from prescription patterns [68]. For instance, proton pump inhibitors might appear beneficial for cardiovascular outcomes in observational data, not due to a therapeutic effect but due to a confounding pattern that varies across institutions, namely co-prescription with aspirin for cardioprotection. CRec separates drug molecular representations into causal therapeutic components and spurious prescription patterns through interventional disentanglement. In an analysis of 47,326 patients with multiple chronic conditions, CRec identified mechanistically synergistic drug combinations that traditional methods missed because they were rarely co-prescribed, revealing genuine therapeutic opportunities obscured by prescribing traditions [68]. The DIR-GNN framework extends invariant learning through environment-based partitioning, explicitly learning which features remain predictive across all environments (causal invariants) versus environment-specific features (spurious correlations), then building predictions based solely on invariant features [69].

### v4

@@ 1 Introduction
In 2019, a widely deployed risk prediction algorithm for the allocation of healthcare resources systematically underestimated the severity of the disease for Black patients by relying on healthcare costs as a proxy for health needs, resulting in discrimination affecting millions of patients annually [1]. That same year, a diabetic retinopathy screening system with specialist-level laboratory performance (>>90% sensitivity) faced critical failure when deployed in community clinics. The model struggled with environmental distribution shift, rejecting 21% of patient images as ungradable due to lighting and protocol differences rarely seen in training data, thereby creating workflow bottlenecks rather than improving care. These failures exemplify the triple crisis of AI in healthcare: models collapse under distribution shift, perpetuate discriminatory patterns embedded in historical data, and struggle to provide reliable mechanistic insight [2, 3]. The paradox of pneumonia risk prediction starkly illustrates this brittleness: models learnt to classify asthma patients as lower risk because they received aggressive early treatment in training hospitals, a spurious correlation that could lead to dangerous under-treatment when deployed elsewhere [4]. The question facing healthcare AI is no longer whether systems can achieve impressive accuracy in retrospective test sets, but whether they can survive clinical deployment across diverse institutions and maintain both effectiveness and equity.
@@
This Perspective examines causal graph neural networks, methods that integrate causal inference principles with graph-based deep learning architectures, as a methodological framework for transforming AI in healthcare from reactive pattern recognition to proactive causal medicine. Unlike previous papers that addressed GNNs in healthcare [34] or causal inference more broadly [14, 35] separately, we focus specifically on their integration, methods in which both graph structure and causal reasoning prove essential for clinical impact. We analyse how causal principles enhance GNN architectures for healthcare tasks, moving beyond correlation to mechanism, and examine current applications that demonstrate clinical utility across psychiatric diagnosis, cancer treatment, physiological monitoring, and drug discovery. We also identify fundamental challenges in scaling these methods to real-world deployment, including computational complexity, validation without ground truth, fairness under distribution shift, and regulatory considerations. Our analysis reveals both transformative promise and substantial barriers: while CIGNNs achieve superior generalisation across healthcare institutions, identify biologically plausible mechanisms, and maintain fairness properties under distribution shift, they face computational requirements precluding real-time application, difficulty validating causal claims from observational data, and gaps between methodological sophistication and clinical interpretability.
@@ 2.1 Why biological systems demand graph-structured causal models
Biological systems exhibit causal relationships organised as networks across multiple scales of organisation. At the molecular level, gene regulatory networks encode transcriptional control in which regulatory proteins causally influence the expression of target genes, with modern single-cell RNA sequencing that captures the expression profiles of 20,000+ genes simultaneously [28]. Protein-protein interaction networks represent physical binding relationships that mediate cellular signalling, with causal perturbations propagating through these networks to generate phenotypic responses [16]. At the tissue level, brain networks comprise hundreds of regions with functional and structural connectivity encoding causal influences between neural populations, where activity in one region can drive or modulate activity in connected regions [40]. At the population level, disease comorbidity networks represent relationships where large-scale data mining of electronic health records and genetic association data are integrated into knowledge graphs to reveal how conditions are linked through shared molecular features [17]. This inherent graph structure makes neural network architectures that explicitly represent and reason about network topology the natural framework for biomedical machine learning.
@@ 2.2 From association to intervention to counterfactuals
Counterfactual reasoning operates at the highest level of Pearl’s hierarchy, addressing patient-specific questions: would this patient have developed complications if we had chosen alternative treatment? Reasoning of this kind is essential for precision medicine, treatment optimisation, and retrospective analysis of clinical decisions [13]. The calculation of counterfactuals requires a three-step process, including abduction (inferring unobserved variables, including patient-specific factors, from observed data), action (modifying the causal model to reflect alternative intervention), and prediction (computing outcomes under the modified model) [41]. The CLEAR framework implements counterfactual generation for graphs through variational autoencoders that learn distributions over causally plausible graph modifications. This could allow, for example, the exploration of alternative biological configurations with some mechanistic validity [51]. In cancer treatment planning, this makes it possible to simulate how different therapeutic targets would affect tumour evolution, identifying vulnerabilities that might not be apparent from observational data alone.
@@ 3.1 Disentangling causal signal from confounding
The DisC (Debiasing via Disentangled Causal Substructure) framework operationalises this separation through a dual-encoder architecture [49]. Rather than treating graphs monolithically, DisC maintains parallel GNN encoders: one trained to extract causal features that consistently predict outcomes across different environments, another capturing environment-specific spurious patterns. The contrastive learning objective enforces the independence between these representations, ensuring that causal and spurious features remain disentangled. For psychiatric diagnosis, this architectural separation was proven to be immensely effective. Traditional GNNs identified widespread hyperconnectivity in the cerebellar regions of depressed patients, a finding inconsistent with the clinical understanding that depression primarily influences affective rather than motor circuits.
@@
The CI-GNN (Causality-Inspired Graph Neural Network) advances this concept through Granger causality-inspired conditional mutual information, providing principled measures of causal strength for graph edges [50]. The model identifies influential subgraphs that demonstrate genuine causal connections through disentangled subgraph-level representation learning. CI-GNN’s causal decomposition correctly identified sparse cerebellar involvement while highlighting strong connections around the left rectus, a region previously identified as critical for antidepressant treatment response prediction [50]. Evaluated on the REST-meta-MDD dataset (n=1,604), this alignment with established neurobiology demonstrates that causal disentanglement provides mechanistically meaningful insights, not merely improved prediction. In autism spectrum disorder evaluations on the ABIDE dataset, CI-GNN revealed tight connections within the somatomotor network consistent with motor impairments characteristic of ASD, findings replicated across independent cohorts and suggesting genuine causal relationships rather than dataset-specific artefacts [50]. The clinical imperative was that clinicians needed to know which connections to target through neurostimulation, pharmacotherapy, or behavioural interventions, not merely that connectivity is abnormal. Although developed for other applications, IGCL-GNN extends these principles through multi-objective optimisation that simultaneously enhances causal features and suppresses environment-specific spurious patterns, potentially being transferrable to help with robust clinical deployment which requires balancing predictive performance on current datasets with generalisation to unseen environments [56].
@@
These findings depend on the chosen parcellation resolution. CI-GNN operates on 116-node brain atlases where each node aggregates thousands of neurons; causal relationships at this mesoscale resolution reflect functional circuit organisation rather than individual neuronal connectivity, with direct implications for which neurostimulation or pharmacological interventions could target the identified causal pathways. Figure 3 illustrates these architectural principles applied to functional neuroimaging data. While the DisC and iVGAE frameworks were originally developed on synthetic benchmark datasets, their conceptual translation here to brain connectivity analysis could demonstrate how causal disentanglement and interventional prediction could address the specific confounding challenges in clinical neuroimaging, separating genuine neural circuit dysfunction from global arousal states, scanner artefacts, or site-specific acquisition protocols.
@@ 3.2 Predicting interventional effects without experimental data
Precision medicine requires not just prediction, but mechanistic explanation, identifying the specific biological features driving a diagnosis. The RC-Explainer (Reinforced Causal Explainer) uses reinforcement learning to discover which edges of a graph to retain to maximise the causal information flow [57]. Rather than computing interventions symbolically, it learns a policy for selecting explanatory subgraphs, with each candidate evaluated not by correlation but by causal impact. In molecular property prediction, this approach identified causal functional groups (e.g., toxicophores) determining mutagenicity, distinguishing active substructures from inert scaffolds, mechanistic insights missed by correlation-based methods that cannot distinguish causal features from spurious environmental associations [57].
For blood pressure monitoring, the CiGNN framework demonstrates the clinical utility of interventional prediction by first using the neural-accelerated causal discovery (FCI algorithm) to identify physiological signals causally related to blood pressure variation, then constructing spatiotemporal GNN architectures guided by the discovered causal structure [48, 58]. This causal feature selection reduces dimensionality from 222 initial wearable features to a sparse set of direct and indirect causal indicators while improving both accuracy and interpretability, identifying, for example, that specific PPG morphological features (e.g., pulse arrival time, waveform characteristics) causally influence blood pressure through arterial stiffness mechanisms.
@@ 3.3 Generating counterfactuals for personalised medicine
Counterfactual reasoning addresses the highest level of Pearl’s causal hierarchy: what would have happened to this patient under alternative treatment? Such patient-specific predictions are essential for precision medicine, treatment optimisation, and retrospective analysis of clinical decisions [13]. The CLEAR (Counterfactual Learning on Graphs) framework approaches counterfactual generation through variational autoencoders that learn distributions over possible graph modifications [51]. Rather than generating arbitrary alternative graphs, CLEAR learns which modifications are causally plausible, which edges could change, maintaining validity relative to the data distribution. In molecular discovery, this leads to exploration of how structural modifications would affect molecular properties, identifying functional groups that drive toxicity or efficacy, vulnerabilities invisible in observational data alone [51].
A challenge in counterfactual reasoning is that interventions can create graph configurations far outside the training distribution. When simulating disruption of a critical hub protein, we can query a configuration likely never appearing naturally, and thus unrepresented in training data. Traditional neural networks can fail potentially under such a distribution shift. For epileptic seizure prediction, dynamic graph models like DIGLN capture the temporal evolution of brain networks [59], future extensions could enable in silico testing of interventions, simulating how disrupting specific neural circuits would affect seizure propagation patterns, guiding neurosurgical planning or neurostimulation protocols.
The CXGNN (Causal Explainer based on GNN-Neural Causal Models) framework provides a rigorous quantification of counterfactual effects through neural implementations of structural causal models [60]. By parameterising each structural equation as a neural network, CXGNN learns arbitrary non-linear causal relationships and stays true to the formal guarantees of SCM-based counterfactual computation. For mild cognitive impairment progression, CXGNN identified that hypertension, arrhythmia, and sex emerged as critical causal factors, with notable sex-specific differences: female patients showed stronger causal links between vascular factors and cognitive decline, while male patients exhibited more pronounced metabolic effects [60, 61]. These mechanistic insights enable personalised risk assessment and targeted early intervention strategies. These counterfactual generation methods form the computational engine of Causal Digital Twins, the ability to simulate patient-specific responses to interventions never administered in reality constitutes the core functionality enabling in silico clinical experimentation.
@@ 3.4 Achieving robustness across heterogeneous clinical settings
While disentanglement methods (Section 3.1) achieve robustness by explicitly separating causal from spurious representations, invariant learning approaches pursue the same objective through a complementary mechanism: optimisation constraints that enforce feature stability across environments without requiring explicit architectural separation. Disentanglement provides interpretable causal-spurious decomposition and invariant learning provides deployment guarantees with potentially simpler architectures. Some methods, such as IGCL-GNN, integrate both principles.
@@
The IGCL-GNN framework relies on invariant causal learning through multi-objective optimisation that enhances causal features and suppresses environment-specific spurious patterns [56]. Though developed for industrial sensor networks, it provides a template for robust clinical deployment requires balancing competing objectives such as maximising predictive performance on current datasets and also ensuring generalisation to unseen environments. In multi-site brain imaging studies, CI-GNN achieved mean accuracy of 0.71–0.73 across 17 acquisition sites compared to 0.64–0.69 for baseline methods, with consistent improvements in 13 of 17 sites despite individual site accuracy ranging from 0.60 to 0.85, demonstrating the value of learning causal invariances rather than site-specific correlations [49, 50]. This stability under distribution shift is not just a technical achievement, but a prerequisite for clinical trust, models that collapse when moved between hospitals cannot be deployed safely.
The CRec (Conditional Causal Representation Learner) framework addresses the robustness of drug recommendation through disentanglement of therapeutic effects from prescription patterns [63]. For example, proton pump inhibitors might appear beneficial for cardiovascular outcomes in observational data, not due to a therapeutic effect but due to a confounding pattern that varies across institutions, namely co-prescription with aspirin for cardioprotection. CRec separates drug molecular representations into causal therapeutic components and spurious prescription patterns through interventional disentanglement. In an analysis of 47,326 patients with multiple chronic conditions, CRec identified mechanistically synergistic drug combinations that traditional methods missed because they were rarely co-prescribed, revealing genuine therapeutic opportunities obscured by prescribing traditions [63]. The DIR-GNN framework extends invariant learning through environment-based partitioning, explicitly learning which features remain predictive across all environments (causal invariants) versus environment-specific features (spurious correlations), then building predictions based solely on invariant features [64].